7. An airplane covered 11 miles of its route while decreasing its altitude by 29,000 feet. Find the slope of the airplane's line of descent. Round to the nearest hundredth. [Hint: 1 mi = 5280 feet.]
no book Thank you very much for all you do to help those who struggle!

Answer by Nate(3500)

This makes a right triangle.
Leg: 29000ft
Leg: 11mi or 58080ft
c^2 = a^2 + b^2
c = sqrt(3373286400 + 841000000)
c = about 64,917.54ft or about 12.29mi